OpenStudy (anonymous):

We can see that 3*4*5*6 = 360 = 19^2-1. Prove that the product of four consecutive integers is always one less than a perfect square.

OpenStudy (anonymous):

(x)(x+1)(x+2)(x+3)=6 x + 11 x^2 + 6 x^3 + x^4 6 x + 11 x^2 + 6 x^3 + x^4=a^2 - 1 a=\[\sqrt{1+6 x+11 x^2+6 x^3+x^4}\] a=\[\left(1+3 x+x^2\right)\]

OpenStudy (anonymous):

2 x 3 x4 x 5=120

OpenStudy (anonymous):

This is really a not good proof but it does find what number have to be squared.

OpenStudy (anonymous):

why dont you like it?

OpenStudy (anonymous):

11^2-1=120

OpenStudy (anonymous):

2*3*4*5=120 which is 121(perfect square -1)

OpenStudy (radar):

Ah so, maybe you come up with something good lol

OpenStudy (anonymous):

because it only find a, not really a definite proof

OpenStudy (radar):

Yeah, some math magic.